### v2

# Disentangling Memory and Reasoning Ability in Large Language Models

Abstract
Large Language Models (LLMs) have demonstrated strong performance in handling complex tasks requiring both extensive knowledge and reasoning abilities. However, the existing LLM inference pipeline operates as an opaque process without explicit separation between knowledge retrieval and reasoning steps, making the model’s decision-making process unclear and disorganized. This ambiguity can lead to issues such as hallucinations and knowledge forgetting, which significantly impact the reliability of LLMs in high-stakes domains. In this paper, we propose a new inference paradigm that decomposes the complex inference process into two distinct and clear actions: (1) memory recall: which retrieves relevant knowledge, and (2) reasoning: which performs logical steps based on the recalled knowledge. To facilitate this decomposition, we introduce two special tokens and , guiding the model to distinguish between steps that require knowledge retrieval and those that involve reasoning. Our experiment results show that this decomposition not only improves model performance but also enhances the interpretability of the inference process, enabling users to identify sources of error and refine model responses effectively. The code is available at https://github.com/MingyuJ666/Disentangling-Memory-and-Reasoning.
1 Introduction
Recent advancements in Large Language Models (LLMs) have showcased their impressive inference capabilities in handling complex natural language tasks that require both extensive knowledge and sophisticated reasoning abilities (OpenAI, 2024; Touvron et al., 2023; Wei et al., 2022a). LLMs have demonstrated the ability to memorize vast amounts of knowledge, and techniques like Chain-of-Thought (CoT) (Wei et al., 2022b), Tree of thoughts (ToT) (Yao et al., 2024) have been developed to further enhance their inference abilities by decomposing complex problems into several simpler, single-step processes. These methods enable LLMs to tackle multi-step inference tasks more effectively by organizing the thought process into discrete, focused actions (Feng et al., 2024; Jin et al., 2024b; Wei et al., 2022b).
However, despite these advancements, existing inference frameworks often operate as an opaque process without explicit separation between knowledge retrieval and reasoning steps. This makes it unclear what specific knowledge the model utilizes and how it performs reasoning, leaving the decision-making process ambiguous. For complex, knowledge-intensive tasks, such as multi-hop inference, LLMs often struggle to effectively leverage their memory for inference (Yang et al., 2023; Jin et al., 2024b; Wang et al., 2024b; Cheng et al., 2024; Liu et al., 2024). Such tasks typically require the ability to recall relevant knowledge for each reasoning step (or "hop") and then perform inference over that recalled memory (Wang et al., 2024c). The lack of structure in the output and effective memory utilization can lead to issues such as hallucinations, where LLMs generate plausible but incorrect information (Xu et al., 2024; Li et al., 2024a), and "forgetting," where relevant information is lost across reasoning steps (Jin et al., 2024b; Chen & Shu, 2023), disrupting the logical flow. These issues compromise the LLM’s accuracy and reliability, posing serious risks in high-stakes applications like healthcare and finance Pham & Vo (2024).
Existing works on enhancing inference in LLMs to address their challenges can be categorized into two approaches: Memory-Based Approaches: These methods focus on improving the recall and utilization of world knowledge stored in the model, such as leveraging Retrieval-Augmented Generation (RAG) (Cai et al., 2019; Chen et al., 2024b). The emphasis is on enabling models to access and use their outside knowledge more effectively. Reasoning-Based Approaches: These techniques aim to improve the reasoning capabilities of models by Chain-of-Thought (CoT) reasoning (Yang et al., 2023; Gao et al., 2024; Yu et al., 2024) or introducing structured guidance in training such as planning tokens (Wang et al., 2023b, 2024b) to organize reasoning into discrete, interpretable steps. These methods enhance the ability of LLMs to handle complex reasoning tasks by embedding structural reasoning mechanisms into their parameters. Despite advancements in both categories, LLMs still struggle with tasks requiring an intricate interplay of memory recall and logical reasoning (Wang et al., 2024c).
In this work, we introduce a new LLM inference paradigm that decomposes the complex inference process into two distinct actions: memory and reasoning. To implement this paradigm, we first construct a new training dataset by leveraging a powerful LLM such as GPT-4o to generate itemized actions for various question-answering tasks, with each action labeled as either reason or memory. Each action is preceded by a special token, either or , which serves as a control signal during training. The second step involves training a customized LLM using this labeled dataset. The inclusion of these two learnable control tokens guides the model to distinguish between recalling relevant knowledge and performing reasoning steps. With these control tokens, we encourage the trained model to first use memory to recall relevant knowledge and then apply reasoning based on that memory to answer the question. This approach not only constrains the model’s behavior but also enhances the interpretability of the inference process.
Our experimental results demonstrate that the proposed decomposition improves performance and enhances the interpretability of the model’s inference process. Specifically, our method achieves accuracy of 78.6% and 78.0% on the StrategyQA dataset (Geva et al., 2021) using Qwen2.5-7B (Yang et al., 2024a) and LLaMA-3.1-8B (Touvron et al., 2023), respectively. These results represent improvements of 1.2% and 1.3% over the planning-token fine-tuned baseline while remaining only 2.2% below GPT-4o’s performance. Remarkably, on the TruthfulQA dataset (Lin et al., 2022), LLaMA-3.1-8B enhanced by our algorithm outperforms GPT-4o with Chain of Thought prompting (85.4%), achieving 86.6% accuracy. On average across three benchmark datasets, our method narrows the performance gap with the top-performing closed-source model, GPT-4o (using CoT prompting), to just 1.9%. Furthermore, our algorithm achieves a disentanglement accuracy of 83.3%, providing insights into the model’s inference. Analyzing the errors made by LLaMA-3.1-8B reveals that most issues stem from reasoning, rather than deficiencies in the knowledge itself. This distinction sheds light on the primary sources of errors in the model’s outputs and enables targeted improvements. Notably, our evaluation approach incorporates a detailed error ratio analysis that distinguishes between memory and reasoning steps, offering a more nuanced understanding of model limitations compared to traditional accuracy metrics. By identifying specific areas for refinement, this analysis supports more effective and informed enhancements to model performance.
Our main contributions are as follows:
New Inference Paradigm for LLMs: We introduce a framework that decomposes inference in LLMs into memory and reason steps, guiding the model to separate knowledge retrieval from logical reasoning, thus enhancing performance and interpretability.
Improve Benchmark Performance: Our model achieves competitive results, surpassing GPT-4o on TruthfulQA and closely matching GPT4-o on StrategyQA and CommonsenseQA, demonstrating the benefits of our approach.
Enhance Interpretability and Control: Our framework enables transparent reasoning with labeled steps for memory and reasoning, allowing precise error analysis and control in high-stakes applications.
2 Method
The workflow of our algorithm is illustrated in Figure 2. Given a question, we first adopt an LLM generation framework, consisting of a knowledge LLM to extract relevant knowledge, followed by a reasoning LLM to derive logical steps. The data we generated for the question are annotated in steps of and . Finally, the questions with annotated steps and answers are then combined to train the model, where the annotations serve as a special trainable token.
2.1 Data Generation
In this section, we introduce an LLM-based framework to generate memory (knowledge in LLM) and reasoning steps, consisting of an inference LLM and a knowledge LLM, as illustrated in Figure 2. First, we use an inference LLM to generate Chain of Thought (CoT) (Wei et al., 2022b) inference steps, prompting it to mark steps that require factual knowledge as and those requiring reasoning as . To improve the quality of the memory steps, we further instruct the inference LLM to rephrase knowledge marked as into questions, emphasizing its factual nature. For the example question in Figure 1, The inference LLM first retrieves relevant knowledge about MMA and Roman Colosseum games, analyzes their relationship and synthesizes this information to form a coherent judgment. By methodically aligning each step with its purpose, the LLM ensures that the conclusion—MMA is not "totally original" from the Colosseum games—reflects well-supported reasoning. Next, a knowledge LLM answers the additional factual questions generated by the inference LLM, such as What is mixed martial arts? and What are the Roman Colosseum games?. The answers to these questions are then substituted into the inference steps. This approach not only effectively decouples reasoning and knowledge to ensure accuracy but also maintains high data quality. This data generation approach is designed to fine-tune LLMs so that LLMs can effectively disentangle knowledge and reasoning during inference.
2.2 Training
In this method, we train a LLM by incorporating intervened reasoning and memory processes as Figure 2, guided by two distinct tokens: , which represents reasoning with knowledge, and , which signifies retrieved knowledge. These tokens are designed to prompt the model to activate specific reason and knowledge abilities, strengthening its reasoning capabilities and ultimately enhancing both interpretability and performance in complex inference tasks.
In training time, each training instance comprises the following components: (1) the question tokens where is the question token length, (2) the step-by-step thinking process consists of intertwined memory and reasoning elements, denoted as and , where each is initiated by a special token followed by a sequence of tokens that represent retrieved factual or stored knowledge: , and is initiated by a special token followed by a sequence of tokens that represent the reasoning process: , and (3) the target answer generated after the completion of the memory retrieval and reasoning processes. The model is trained in a standard autoregressive manner using LoRA fine-tuning and the and are trainable out of vocabulary tokens.
By structuring the input in this manner, the model learns to process and distinguish between retrieved knowledge and the reasoning steps required to generate the final answer. The inclusion of the and tokens facilitates the disentanglement of memory retrieval and reasoning processes, thereby enhancing the model’s ability to produce coherent and accurate responses.
3 Experiment
3.1 Experiment Setup
Models.
In our experiments, we use LLaMA-2-7B-chat-hf (Touvron et al., 2023), LLaMA-3.1-8B-Instruct (Dubey et al., 2024), and Qwen2.5-7B-Instruct (Yang et al., 2024a) as target models for training. GPT-4o serves as the inference and knowledge LLM to generate training data, while GPT-4o-mini is employed as the evaluator.
Datasets.
Our experiments are conducted on three datasets. StrategyQA (Geva et al., 2021) is a question-answering benchmark comprising 2,780 examples that require inference, with each example containing a strategy question, a Chain-of-Thought (CoT) path, and supporting evidence. CommonsenseQA (Talmor et al., 2019) is a question-answering dataset with 12,102 questions, each necessitating commonsense knowledge to identify the correct answer among four distractors. TruthfulQA (Lin et al., 2022) serves as a benchmark for evaluating the truthfulness of language model responses, encompassing 817 questions across 38 categories. For TruthfulQA, we use the mc1_targets, which contains single-choice questions with 4-5 answer choices per question. We transformed this dataset into a single-choice format to facilitate training and testing with answer options labeled as A-E. In the original dataset, the correct answer was always the first option. We shuffled the labels to prevent the model from learning shortcuts. StrategyQA and CommonsenseQA datasets have predefined train and test splits for mc1_targets in TruthfulQA, which we divided into an 8:2 ratio for the training and testing sets.
Baselines.
Five main methods are being compared: Zero-shot, CoT (Chain-of-Thought), LoRA, and Planning-token. In our experiments, we adopt zero shot (Just input the question) and Chain of Thought (CoT) prompting as our vanilla baseline for reasoning. For the fine-tuned baseline, we use int8_training to save memory and we choose LoRA fine-tuning (Hu et al., 2021) and Planning Tokens (LoRA+Prompt Tuning) (Wang et al., 2023b) to train and test on these three datasets to facilitate a comparative evaluation with our proposed approach.
Evaluation Metric.
We use accuracy (acc) to measure the model’s performance on multiple-choice questions.
3.2 Main Results
Our algorithm consistently achieves superior performance across multiple benchmarks compared to existing approaches. On StrategyQA, LLaMA-3.1-8B fine-tuned by our method achieves 78.0%, outperforming CoT (69.4%) and Planning-token (76.7%), while LLaMA-2-7B fine-tuned with our method achieves 70.6%, compared to LoRA (61.2%) and Planning-token (63.5%). Similarly, in CommonsenseQA, "Ours" with LLaMA-3.1-8B scores 82.3%, surpassing CoT (70.6%) and Planning-token (76.9%), while LLaMA-2-7B with our method achieves 71.1%, outperforming LoRA (65.4%). A significant breakthrough is observed on the TruthfulQA dataset, where LLaMA-3.1-8B enhanced by our method achieves 86.6%, outperforming GPT-4o by 1.8% in zero-shot setting and 1.2% in CoT setting, and also surpassing Planning-token (82.5%) and LoRA (79.8%) in fine-tuned settings. Notably, on StrategyQA, Qwen2.5-7B fine-tuned with our method achieves 78.6%, exceeding Planning-token performance (77.4%) and LoRA (69.1%). On the TruthfulQA dataset, Qwen2.5-7B struggles with CoT settings, achieving only 56.7%, which underscores the instability of instruction-tuned models like Qwen2.5-7B on certain datasets. These results emphasize the robustness and adaptability of our approach in handling inference-intensive tasks across diverse datasets and models. However, adding a CoT inference step can sometimes decrease performance, as also reported by Sprague et al. (2024). The unstable training of instruction-tuned models like Qwen2.5-7B further highlights the limitations of existing methods compared to our approach.
| Methods | Models | StrategyQA | CommonSenseQA | TruthfulQA | Average |
| Vanilla | |||||
| LLaMA-2-7B | 0.607 | 0.523 | 0.262 | 0.464 | |
| Zero-shot | LLaMA-2-13B | 0.613 | 0.530 | 0.378 | 0.507 |
| LLaMA-3.1-8B | 0.659 | 0.635 | 0.616 | 0.637 | |
| LLaMA 3.1-70B | 0.796 | 0.765 | 0.793 | 0.785 | |
| Qwen 2.5-7B | 0.640 | 0.789 | 0.726 | 0.718 | |
| GPT-4o | 0.699 | 0.834 | 0.848 | 0.794 | |
| LLaMA-2-7B | - | - | - | - | |
| CoT | LLaMA-2-13B | 0.560 | 0.482 | 0.390 | 0.477 |
| LLaMA-3.1-8B | 0.694 | 0.706 | 0.506 | 0.635 | |
| LLaMA 3.1-70B | 0.822 | 0.815 | 0.762 | 0.800 | |
| Qwen 2.5-7B | 0.696 | 0.784 | 0.567 | 0.682 | |
| GPT-4o | 0.808 | 0.865 | 0.854 | 0.842 | |
| Fine-tuned | |||||
| LLaMA-2-7B | 0.612 | 0.641 | 0.767 | 0.673 | |
| LLaMA-2-13B | 0.696 | - | - | 0.696 | |
| LoRA | LLaMA-3.1-8B | 0.701 | 0.754 | 0.798 | 0.737 |
| Qwen 2.5-7B | 0.691 | 0.775 | 0.725 | 0.730 | |
| LLaMA-2-7B | 0.635 | 0.654 | 0.770 | 0.686 | |
| LLaMA-2-13B | 0.715 | - | - | 0.715 | |
| Planning-token | LLaMA-3.1-8B | 0.767 | 0.769 | 0.825 | 0.787 |
| Qwen 2.5-7B | 0.774 | 0.801 | 0.762 | 0.779 | |
| LLaMA-2-7B | 0.706 | 0.711 | 0.786 | 0.734 | |
| LLaMA-2-13B | 0.739 | - | - | 0.739 | |
| Ours | LLaMA-3.1-8B | 0.780 | 0.823 | 0.866 | 0.823 |
| Qwen 2.5-7B | 0.786 | 0.832 | 0.812 | 0.810 | |
3.3 Ablation Study
3.3.1 The contribution of Memory and Reason token?
The Ablation Experiment presents an ablation study comparing the performance of two versions of the LLaMA model ( LLaMA-2-7B and LLaMA-3.1-8B) across three benchmarks: StrategyQA, CommonsenseQA, and TruthfulQA. The study examines the impact of using specific tokens ("Memory and Reason" vs. "Random") on the model’s performance. During training, we shuffled the allocation of and tokens and then observed the effects on training and testing performance. As expected, the overall performance declined, but the rate of decline varied, demonstrating the superiority of our approach in disentangling reason and memory.
3.3.2 The impact of the number of special tokens
In our training setup, in addition to the two token types, and , we include a parameter representing the number of special tokens preceding each sentence. A sentence might include three tokens or four like the example above, with the same variation applied to memory tokens. Our experiments indicate that model performance improves with around four to six special tokens, likely because more tokens can lead to better performance for the model during training (Levy et al., 2024). As shown in table 4 and table 4, we selected two models ( LLaMA-2-7B, LLaMA-3.1-8B) to illustrate their performance across different numbers of special tokens.
| StrategyQA | CommonsenseQA | TruthfulQA | Average | |
| LLaMA-2-7B | ||||
| w Memory and Reason token | 0.706 | 0.711 | 0.786 | 0.734 |
| w Random token | 0.644 | 0.651 | 0.708 | 0.668 |
| LLaMA-3.1-8B | ||||
| w Memory and Reason token | 0.780 | 0.823 | 0.866 | 0.823 |
| w Random token | 0.759 | 0.795 | 0.840 | 0.798 |
| Model Name | Number of Tokens | |||
| 0 | 2 | 4 | 6 | |
| LLaMA-2-7B | 0.682 | 0.704 | 0.710 | 0.711 |
| LLaMA-3.1-8B | 0.783 | 0.816 | 0.823 | 0.820 |
| Qwen2.5-7B | 0.799 | 0.813 | 0.832 | 0.813 |
| Model Name | Number of Tokens | |||
| 0 | 2 | 4 | 6 | |
| LLaMA-2-7B | 0.701 | 0.762 | 0.768 | 0.786 |
| LLaMA-3.1-8B | 0.826 | 0.865 | 0.866 | 0.859 |
| Qwen2.5-7B | 0.807 | 0.756 | 0.812 | 0.799 |
Another important issue is knowledge distillation. We need to ensure that the model’s improvement is not due to knowledge distillation from the GPT-4 framework. By using the same inference steps, we compared the results of standard training with 0 reason and memory tokens and found that adding these tokens significantly improves performance. This indirectly confirms that the model’s enhancement comes from algorithmic improvements rather than knowledge distillation.
3.4 Case Study
3.4.1 Sample Analysis
To validate the decoupling effect on memory and reasoning, we configure and employ GPT-4o-mini as an evaluator, assessing whether steps labeled as "memory" entail factual knowledge and those labeled as "reasoning" represent reasoning processes on StrategyQA, CommonSenseQA, and TruthfulQA. We use a structured, directive, one-shot Chain of Thought (CoT) prompting method as the baseline to compare with our method. The prompt setup is detailed in Appendix F.
In this CoT approach, directive prompting with explicitly instructs the model to distinguish memory information and reasoning steps. The one-shot example provides a structured format, demonstrating how memory (e.g., ) and reasoning parts (e.g., ) can be organized separately in answer generation (e.g., ). This structure guides the model to produce answers annotated as either memory or reasoning , enhancing interpretability by separating factual knowledge from reasoning processes.
From Figure 3 and Table 7, on the StrategyQA dataset, our method achieves an accuracy of 78.0% on LLaMA-3.1-8B, outperforming the One-shot CoT baseline by 20%, our approach achieves higher accuracy in decoupling memory (94% vs. 93%) and reasoning (71% vs. 67%), demonstrating effective decoupling between these two components in multi-steps inference. On the CommonsenseQA dataset, our method achieves an accuracy of 82.3% on LLaMA-3.1-8B, exceeding the One-shot CoT baseline by 26.3%. The results highlight that our approach consistently outperforms the baseline in decoupling memory (91% vs. 83%) and reasoning (78% vs. 74%), demonstrating robust performance in commonsense inference tasks. On the TruthfulQA dataset, our method achieves an accuracy of 86.6% on LLaMA-3.1-8B, surpassing the One-shot CoT baseline by 32.6%. The results further illustrate that our approach achieves superior accuracy in decoupling memory (89% vs. 88%) and reasoning (74% vs. 62%), highlighting its effectiveness in factual reasoning. Additionally, Table 7 shows that both LLaMA-3.1-8B and LLaMA-2-7B maintain consistent distributions of memory and reasoning across all datasets. This reflects the stability and generalizability of our decoupling mechanism, ensuring its applicability to diverse reasoning tasks.
| Method | Ratio |
| StrategyQA | |
| LLaMA-2-7B | 4:5 |
| LLaMA-3.1-8B | 1:1 |
| CommonsenseQA | |
| LLaMA-2-7B | 1:5 |
| LLaMA-3.1-8B | 1:5 |
| TruthfulQA | |
| LLaMA-2-7B | 3:7 |
| LLaMA-3.1-8B | 1:2 |
| Method | Accuracy (% ) |
| StrategyQA | |
| One-shot CoT | 58.0 |
| Ours | 78.0 |
| CommonsenseQA | |
| One-shot CoT | 56.0 |
| Ours | 82.3 |
| TruthfulQA | |
| One-shot CoT | 54.0 |
| Ours | 86.6 |
| Error Type | Proportion |
| StrategyQA | |
| Memory | 1.7 |
| Reason | 98.3 |
| CommonsenseQA | |
| Memory | 21.6 |
| Reason | 78.4 |
| TruthfulQA | |
| Memory | 21.1 |
| Reason | 78.9 |
3.4.2 Attention Analysis
In the case study, we have found that the and do an important job in our LLM’s reasoning. Although using raw attention weights to interpret token importance can be somewhat controversial, attention patterns still provide valuable insights into how Transformers operate (Abnar & Zuidema, 2020). This heatmap, as Figure 4 shows that the model focuses intensely on specialized tokens throughout the CoT steps. These tokens received higher attention weights than regular tokens, suggesting they may play a more significant role in leading knowledge and reasoning content generation. This observation aligns with the main findings presented in the previous case study.
We input two sentences(which can be found in the appendix) into our fine-tuned LLaMA-3.1-8B model, resulting in a large attention heatmap. We then segmented the entire attention map according to the steps in the CoT, producing the two smaller maps below, as the large map was too cumbersome to analyze. Other attention maps can be found in the appendix. We observed that and tokens had relatively high attention values, indicating that the model places greater emphasis on this content, which indirectly demonstrates the effectiveness of our algorithm.
3.4.3 Error Analysis
We analyzed all incorrect results generated by our fine-tuned LLaMA-3.1-8B model to identify whether the errors originated from memory or reasoning issues, utilizing GPT-4o to categorize the source of each error across StrategyQA, CommonsenseQA, and TruthfulQA benchmarks. As shown in Table 7, 98.3% of the errors in StrategyQA were attributed to reasoning, with only 1.7% due to memory issues, indicating reasoning as the dominant challenge. Similarly, in CommonsenseQA, 78.4% of errors stemmed from reasoning, while 21.6% were caused by memory failures; in TruthfulQA, the trend persisted, with 78.9% of errors linked to reasoning and 21.1% to memory. These results demonstrate that reasoning-related errors consistently account for over 75% of total mistakes across benchmarks, underscoring that while the model successfully utilizes knowledge, it requires significant improvements in reasoning capabilities, pointing to an important direction for future research.
For Example, the question 5 above emphasizes the role of the revolving door as a security measure, so the correct answer should be somewhat unexpected. Options B, C, and D represent typical uses of revolving doors for managing two-way traffic flow. Only in banks does a revolving door serve as a security measure. The correct answer is likely A. bank, as banks use revolving doors not only for easy access but also as a security measure to control entry and exit. The model’s knowledge is accurate, but this nuance was missed during reasoning.
4 Related Work
Parametric Memory in LLMs.
During pre-training, Large Language Models capture a large amount of knowledge in models’ parameters, known as parametric memory (Jin et al., 2024a). Previous research extensively explores the mechanism of inference with parametric memory, they observe that models can well adopt memory for simple tasks but struggle for complex inference, e.g., multi-hop inference (Li et al., 2024b; Yang et al., 2024b; Wang et al., 2024a). Others reveal the challenges in the leverage of parametric knowledge, particularly when dealing with long-tail facts (facts associated with less common entities) or when the knowledge is rare (Wang et al., 2023a; Allen-Zhu & Li, 2023; Cheng et al., 2024). These studies primarily focus on the analysis of model behavior. While valuable, they do not address how to better elicit the parametric knowledge for inference. In this work, we explore how to boost LLMs’ leverage of their parametric knowledge for complex inference.
Reasoning with LLMs.
Recent research on enhancing large language models’ inference capabilities can be broadly categorized into prompt-based and tuning-based approaches. Prompt-based methods strategically guide reasoning processes. Chain-of-Thought (CoT) prompting (Wei et al., 2022b) and its derivatives (Zhao et al., 2024; Zhou et al., 2023; Chen et al., 2024a; Hu et al., 2023; Jin et al., 2024b) decompose complex tasks into sequential steps, improving transparency and decision-making. Others like Tree-of-Thoughts (ToT) (Yao et al., 2023) and Graph-of-Thoughts (GoT) (Besta et al., 2024) further utilize hierarchical and network-based inference to cover larger searching space. These strategies design a framework where LLMs can elicit the parametric memory relevant to the task. However, in these methods, the models might not know when to reason or use their memory. Tuning-based methods introduce trainable tokens for structured CoT steps, facilitating reasoning and utilization of memory Wang et al. (2023b); Goyal et al. (2023); Colon-Hernandez et al. (2024) Despite the effectiveness, these methods intertwine reasoning and memory usage, which may limit the full potential of the models. In contrast, our approach aims to decouple memory and reasoning within the CoT process by introducing various special tokens, enabling the model to leverage its memory more effectively.
5 Conclusion
In this work, we introduced a novel framework for training an autoregressive language model to distinguish between reasoning and memory processes. By utilizing two specific tokens— for retrieved factual knowledge and for reasoning steps—the model can effectively disentangle knowledge retrieval from logical inference, enhancing interpretability and performance on complex reasoning tasks. The structured input encoding enables the model to maintain a clear boundary between memory and reasoning, improving its generalization capabilities for queries requiring a blend of factual knowledge and multi-step reasoning. Through this approach, the model not only achieves more accurate answers but also provides interpretable outputs that clarify the reasoning process step-by-step, a feature critical for tasks in fields requiring transparency and accountability in AI-generated reasoning. Despite the promising results, our approach has certain limitations: As the number of reasoning steps grows, the input sequence length can become substantially large, especially for complex questions requiring multiple reasoning steps. This may lead to increased computational costs and memory limitations, especially when training large models. The model’s interpretability and performance improvements are closely tied to the effectiveness of the and annotations. Incorrect tagging or inconsistent use of these tokens during training could result in ambiguous or incorrect responses during inference.
6 Limitation
The proposed decomposition framework offers a promising approach to disentangling memory recall and reasoning in large language models. However, like any novel framework, its limitations should be acknowledged for future improvement. One limitation is the framework’s reliance on the quality and breadth of training data for memory recall. While this is a common issue for most machine learning models, missing or underrepresented knowledge in the training data could affect the accuracy and completeness of retrieval, particularly in niche domains. However, this limitation can be mitigated by incorporating dynamic updating mechanisms or integrating external knowledge bases to fill in gaps.
Introducing special tokens, such as and , while useful for distinguishing between memory retrieval and reasoning tasks, adds some complexity to the tokenization process. This complexity may require additional task-specific tuning and adaptation to different LLM architectures or language domains. Nevertheless, the token-based approach provides a clear scaffold for interpretability, which could offset the initial complexity by enhancing model transparency. The framework also faces challenges with tasks that require highly complex reasoning, such as multi-hop or deeply nested logic, which may not be easily separable into discrete memory retrieval and reasoning steps. While this suggests the need for further refinement in handling complex chains of thought, the framework still performs well for a wide range of tasks, demonstrating its robustness in standard reasoning scenarios.
Appendix A Appendix
SUMMARY OF THE APPENDIX
This appendix contains additional details for the “Disentangling Memory and Reasoning Ability in Large Language Models”. The appendix is organized as follows:
Appendix B Data Generation
B.1 Implement Details
We developed an LLM-based data generation framework based on GPT-4o to generate high-quality training data for memory and reasoning steps. This framework includes two LLMs: an inference LLM and a knowledge LLM. The inference LLM is responsible for generating Chain-of-Thought (CoT) processes, decoupling memory from reasoning, and assigning labels to each sub-step by marking those requiring factual knowledge as [memory] and those requiring reasoning as [reason]. The prompt of the inference agent is shown as prompt 2 below. The Knowledge LLM retrieves the necessary knowledge for [memory] steps based on questions provided by the inference LLM. The prompt of the Knowledge LMM is shown as prompt 1 below. We use these two LLMs to ensure the independence of memory and reasoning within the CoT, providing high-quality data for subsequent training. Figure 6 is the prompt configuration for a knowledge LLM. The questions corresponds to the Knowledge base content in the inference LLM and is used to supply accurate factual information for steps labeled as <memory>.
The prompt configuration for inference LLM in a zero-shot setting is shown in Figure 7. The Step name refers to the specific name of each step in the Chain of Thought (CoT) process. The Requirement column labels whether each step pertains to <memory> or <reason>. The Knowledge based is used to provide questions related to factual knowledge in <memory> steps, while the Content focuses on designing to outline the reasoning process for <reason> steps. This structure facilitates a clear distinction between memory retrieval and reasoning tasks, enhancing the model’s capability to execute complex sequences in a zero-shot environment.
B.2 Example
In this study, we leverage both the inference LLM and knowledge LLM based on GPT-4o to generate a dataset. Specifically, we use the StrategyQA dataset as the source for our generation. The StrategyQA dataset is a question-answering dataset designed to evaluate models’ multi-hop inference abilities. It includes questions that require strategic thinking and often demand more than one piece of information to answer correctly.
Figure 8 shows enhanced data generated by our data generation agent from the StrategyQA dataset. The image shows an example question about the relationship between Mixed Martial Arts and the origins of Roman Colosseum games. The answering process is broken down into several steps, each labeled as either [Memory] or [Reason] to indicate the type of step. This approach helps differentiate between pure knowledge retrieval and logical reasoning steps, providing more granular training data for models to improve their accuracy and interpretability in answering complex questions.
Appendix C Preliminary
C.1 Experiment
In a preliminary experiment, we analyzed the training and test sets of StrategyQA, TruthfulQA, and CommonsenseQA to evaluate the overlap in knowledge between them. This assessment was crucial to ensure that our model’s performance improvement was due to our advanced algorithm, rather than simply distilling knowledge from GPT-4o. For our synthetic training set, we extracted sentences following each token to create a reference set. We then prompted our fine-tuned LLaMA3.1-8B model to generate outputs using the test set, collecting sentences following the tokens in these outputs to form a separate set. Our validation method involves setting a threshold on cosine similarity and assessing Jaccard similarity based on this threshold. Specifically, as illustrated in Figure 10, we define two knowledge after tokens as overlapping if the cosine similarity of their embeddings exceeds 0.2. Based on this criterion, the Jaccard similarity for all datasets are smaller than 10%, which is a low value indicating a low degree of overlap and demonstrates that our model’s performance is not merely the result of knowledge distillation.
C.2 Example
A value greater than or equal to 0.2 indicates that the two contents are very unrelated, like example 9.
Appendix D Experiment Details
D.1 Dataset
StrategyQA (Geva et al., 2021)
StrategyQA is a challenging question-answering benchmark that focuses on implicit, multi-step reasoning. Unlike conventional multi-hop datasets where questions explicitly outline the steps needed to reach an answer, StrategyQA requires models to infer these reasoning steps. Each question in StrategyQA is crafted to be implicit and short, with Boolean ("Yes" or "No") answers, requiring logical deductions based on general knowledge. For example, answering a question like “Did Aristotle use a laptop?” involves reasoning about the historical timeline of both Aristotle’s life and the invention of laptops. The StrategyQA dataset includes a total of 2,780 verified questions. The training set comprises 1,600 questions which are used for fine-tuning, and the validation (test) set contains 690 questions, which are used for the validation of each baseline and our method in our experiment.
CommonsenseQA (Talmor et al., 2019)
CommonsenseQA is a multiple-choice dataset with 12,247 questions aimed at testing AI on commonsense reasoning using the ConceptNet knowledge graph. Each question has one correct answer and four distractors, requiring models to understand relations like causality and spatial proximity. While humans achieve 88.9% accuracy, advanced models like BERT-Large reach only 55.9%, underscoring the challenge of commonsense reasoning in AI. The training set comprises 9,740 questions, the validation set contains 1,220 questions, which are used for fine-tuning in our experiment, and the tests set includes 1,140 questions, which are used for the validation of each baseline and our method in our experiment.
TruthfulQA (Lin et al., 2022)
TruthfulQA is a benchmark of 817 questions designed to test language models’ truthfulness by prompting common misconceptions across topics like health and law. Models like GPT-3 and GPT-2 often generate false answers that mirror human misunderstandings, with larger models frequently performing worse (58% truthfulness for GPT-3) compared to 94% for humans. The benchmark reveals that scaling up model size alone does not enhance truthfulness, highlighting the need for targeted fine-tuning to reduce imitative falsehoods. In our experiments, we split the dataset into training and testing sets in an 8:2 ratio. Since the original dataset contained only single-choice questions with all answers marked as A, we randomly shuffled the answer options for one question to ensure effective fine-tuning performance on the training set.
D.2 Evaluation Metric
To mitigate the inherent output instability of LLMs in both CoT and Zero-shot settings, we found that conventional answer-matching techniques, such as regular expression-based methods, may not reliably capture the precise answers required. Consequently, we adopted GPT-4o-mini as an evaluation tool to compute the LLM performance across multiple datasets (Cao et al., 2024). This approach enables a more nuanced assessment of LLM outputs, given the limitations of regular matching techniques under these settings. The detailed prompt used for evaluation is shown in Prompt 11 below.
Appendix E Training Details
E.1 Training Configuration
All the experiments for fine-tuning are run on an NVIDIA RTX 6000 Ada Generation GPU. Our experiments found that the optimal configuration for learning out-of-vocabulary (OOV) tokens is with N_PREFIX=3 and N_SPECIAL=4. We generally use a learning rate of 2e-4 with –warmup_steps 1000, –lr_scheduler_type "cosine", and –optim "adamw_torch", along with gradient_accumulation_steps=16. Additionally, we employed int8 training to ensure that the model could be trained on a single GPU. Additionally, We provided detailed parameter configurations as below:
Here is the detailed training configuration:
E.2 Training Process
We monitored the training process of our method across all models and datasets, recording test set accuracy changes every 10 steps, as illustrated in Figure 15.
E.3 Example
In Figure 16, we present an correct example of fine-tuning LLaMa-3.1-8B using our proposed algorithm. The results clearly demonstrate that our method effectively decouples factual knowledge and reasoning steps from inference.
Appendix F Case Study
F.1 Sample Analysis
For sample analysis, to highlight the decoupling effectiveness, reasoning capability, and interpretability of our approach, we set One-shot Chain-of-Thought (CoT) reasoning as the baseline for this evaluation (see details in Figure). We leverage GPT-4o-mini as an evaluator with the prompt configuration provided below to assess the decoupling effectiveness of LLaMA-3.1-8B in separating memory from reasoning processes.
For sample analysis with GPT-4o-mini as evaluator shown in Figure 17 and the one-shot CoT as baseline, we use a sample generated by our data method as an in-context learning example for the one-shot CoT baseline configuration, shown in Figure 22.
F.2 Error Analysis
To ensure accuracy in error detection, we use GPT-4o as an evaluator to assess whether the error occurs in the <memory> or <reasoning> step, based on the correct answer and the provided reasoning process. The prompt configuration of GPT-4o-mini is shown in Figure 18.
F.3 More Attention Maps
In Figure 20, we have shown more examples of More Attention Maps on StrategyQA for LLaMA3.1-8B and LLaMA2-7B. The prompt for the attention map is in Figure 19.
F.4 More Analysis Samples
To validate the decoupling effect on memory and reasoning, we additionally evaluated the performance of our method on LLaMA-2-7B in comparison with one-shot CoT. As shown in Figure 21, our algorithm outperforms one-shot CoT in terms of the decoupling effect, demonstrating the effectiveness of our approach on LLaMA2-7B.
Appendix G Future Work and Limitation
G.1 Future Work
Dynamic Memory Updating.
Future research could explore mechanisms for dynamically updating the model’s memory, allowing it to incorporate new information without extensive retraining. This would help the model stay current and relevant, especially for knowledge that frequently changes.
Adaptive Reasoning Steps.
Developing methods that enable the model to adaptively select the number of reasoning steps based on task complexity would improve both performance and efficiency. This could involve learning when to retrieve memory and when to directly reason, optimizing the inference process.
Interpretable Error Analysis Tools.
Building on the interpretability gains of the proposed framework, future work could focus on developing error analysis tools that make it easier for users to trace specific failures to either memory recall or reasoning steps, aiding in systematic model improvement.
Cross-Domain Generalization.
Extending the proposed method to domains beyond language (e.g., multimodal tasks) could be an interesting direction. By testing and adapting this decomposition in fields such as vision-language tasks, researchers could evaluate its utility in more complex, real-world applications.
User-Guided Memory and Reasoning.
Investigating ways for users to guide or interact with the model’s memory retrieval and reasoning steps, perhaps through feedback loops, could improve user control and trust in model outputs, especially in high-stakes applications.
G.2 Limitation
Dependency on Training Data.
The proposed decomposition framework relies heavily on the quality and breadth of training data for the memory recall process. If certain knowledge is missing or inadequately represented in the training data, the model may still struggle with knowledge retrieval, potentially leading to inaccurate or incomplete responses.
Token Utilization Complexity.
The introduction of special tokens, such as and , while useful, may add complexity to the tokenization process and necessitate further tuning for various tasks. This can make the framework less straightforward to apply across different LLM architectures or language domains.
Performance in Highly Complex Reasoning Tasks.
While the decomposition approach shows promise in improving reasoning interpretability and accuracy, it may still struggle with tasks requiring multi-hop or deeply nested reasoning steps. Complex chains of reasoning may not be easily separated into discrete memory retrieval and reasoning actions.
Computation Overhead.
The process of decomposing memory recall and reasoning steps can increase computation time due to the additional need for retrieval-based processing. This can be a limitation for real-time applications or systems requiring rapid inference.